adam is 26 years younger than his mother. In 4 years his mother will be 3 times Adam's age. How old is Adam now?
Answer by ewatrrr(24785)

x Adam's age
x+26 + 4 = 3(x+4)
x + 30 = 3x + 12
18 = 2x
9 = x, Adam's age. HIs mother is 35
and..Checking..in 4 years
39 = 3*13